OpenStudy (anonymous):

i have a square that has a diagonal measuring 122√ cm long. i cannot find the length of the sides. can someone help.

OpenStudy (jack1):

h^2 = a^2 + b^2 h^2 = 2a^2

OpenStudy (anonymous):

o.o

OpenStudy (jack1):



OpenStudy (anonymous):

O.O

OpenStudy (jack1):



OpenStudy (anonymous):

11

OpenStudy (jack1):

h^2 = 122 = 2a^2 so a^2 = 122/2 = 61 so a = sqrt 61 = like 8ish...?

OpenStudy (jack1):

a = 7.8103

OpenStudy (jack1):

háll good...?

OpenStudy (anonymous):

ok so y didnt u have to sqr the 122?

OpenStudy (anonymous):

srry i meant *find sqr root*

OpenStudy (jack1):

h = sqrt 122 so h^2 = 122 = 2a^2

OpenStudy (anonymous):

ohhhh okay thanks man i appreciate it

OpenStudy (jack1):

totes good, welcomes